# When F1 Fails: Granularity-Aware Evaluation for Dialogue Topic Segmentation

### v2

@@ Abstract
Dialogue topic segmentation supports summarization, retrieval, memory management, and conversational continuity. Despite decades of prior work, evaluation practice in dialogue topic segmentation remains dominated by strict boundary matching and F1-based metrics, even as modern large language model (LLM) based conversational systems increasingly rely on segmentation to manage conversation history beyond the model’s fixed context window, where unstructured context accumulation degrades efficiency and coherence.
This paper introduces an evaluation perspective for dialogue topic segmentation that reports boundary density and segment alignment diagnostics (purity and coverage) alongside window-tolerant F1 (W-F1). By explicitly controlling boundary density through separating boundary scoring from boundary selection, we evaluate segmentation quality across density regimes rather than at a single operating point. Through cross-dataset empirical evaluation, we show that reported performance differences across dialogue segmentation benchmarks are often not driven by boundary placement quality alone; under common evaluation protocols, annotation granularity mismatch and sparse boundary labels can dominate these differences.
We evaluate multiple, structurally distinct dialogue segmentation strategies across eight dialogue datasets spanning task-oriented, open-domain, meeting-style, and synthetic interactions. Across these settings, we observe that boundary-based metrics are strongly coupled to boundary density, with threshold sweeps producing larger W-F1 changes than switching between structurally distinct methods. These findings support viewing topic segmentation as a granularity selection problem rather than as the prediction of a single correct boundary set, and motivate separating boundary scoring from boundary selection as an algorithmic formulation for analyzing and tuning segmentation behavior under varying annotation granularities.
@@ 1 Introduction
Although we do not introduce a new neural architecture for boundary scoring, the proposed separation of boundary scoring and boundary selection motivates alternative ways of reasoning about the construction, tuning, and evaluation of segmentation models under varying annotation granularities.
@@ 1.1 Motivation: Conversational Memory in AI Chat Systems
In this setting, topic segmentation functions as a memory control mechanism rather than a purely linguistic task. Poor segmentation can cause relevant context to be discarded prematurely or irrelevant context to be retained, leading to degraded long-horizon coherence and unstable behavior across turns (e.g., [sclar-etal-2024-quantifying, razavi-etal-2025-promptset]). The lack of robust long-horizon memory remains an open challenge for current LLM-based systems, particularly in conversations spanning multiple sessions, motivating careful attention to segmentation decisions and their evaluation.
@@ 2.1 Boundary-Based Evaluation Metrics
Topic segmentation evaluation has historically centered on boundary alignment. Early metrics such as $P_{k}$ beeferman1999statistical penalize segmentations where predicted and reference boundaries fall in different segments under a sliding window. WindowDiff pevzner2002critique refines this approach to address known biases in $P_{k}$. Both metrics are typically computed against a single reference segmentation and measure deviation from that reference, even though multiple segmentations (and granularities) can be defensible [fournier-inkpen-2012-segmentation, fournier-2013-evaluating]. More recent dialogue segmentation work has largely adopted precision, recall, and exact-match $F_{1}$ computed over boundary positions liu2024dialogseg. These boundary-only scores likewise treat the available annotation as the evaluation target; when annotation granularity varies across datasets or differs from system behavior, they can yield misleading comparisons.
@@
A substantial line of work in text segmentation evaluation argues that (1) strict boundary matching over-penalizes near misses, and (2) window-based metrics such as $P_{k}$ and WindowDiff introduce task- and parameter-dependent biases, motivating edit-distance and confusion-matrix-based alternatives. fournier-inkpen-2012-segmentation propose segmentation similarity ($S$), an edit-distance–based similarity that is symmetric (does not privilege a single “true” reference) and supports evaluation against multiply-coded corpora. fournier-2013-evaluating extend this direction with boundary edit distance and boundary similarity ($B$), yielding a principled way to obtain boundary-level IR-style quantities (precision/recall/$F$) while awarding graded credit for near misses. Relatedly, scaiano-inkpen-2012-getting derive a boundary confusion matrix from WindowDiff-style windows (WinPR), separating false positives from false negatives while retaining tunable near-miss sensitivity. For corpora with multiple annotators or where segmentation is treated as a unitizing task, agreement measures that jointly align and score units (e.g., $\gamma$) provide an alternative to single-reference evaluation [mathet2015gamma].
@@ 2.2 Dialogue Segmentation Datasets
Dialogue topic segmentation introduces complexity beyond written text due to turn-taking, discourse phenomena, and gradual topic transitions. Several benchmarks have emerged with distinct annotation philosophies. DialSeg711 dialseg711 and SuperDialseg superdialseg2023 provide densely annotated dialogue corpora, while TIAGE tiage annotates topic boundaries in task-oriented dialogue with sparser labels tied to task structure. MultiWOZ budzianowski2018multiwoz provides domain annotations that can be interpreted as coarse topic boundaries. Open-domain and semi-structured datasets such as DailyDialog li2020dailydialog, Taskmaster byrne2019taskmaster, Topical-Chat topicalchat, and QMSum zhong2021qmsum reflect further variation in how topic-like structure is operationalized, including domain/section labels or derived boundary heuristics rather than explicit topic-boundary annotation. This diversity in annotation intent produces systematic differences in boundary density across datasets—differences that standard boundary scores do not distinguish from detection quality.
@@ 4 Datasets
We evaluated multiple, structurally distinct dialogue segmentation strategies across eight dialogue datasets spanning diverse dialogue regimes. The purpose of cross-dataset evaluation here is not only to compare performance, but to expose how annotation philosophy and boundary density interact with evaluation metrics.
Task-oriented dialogue datasets include SuperDialseg superdialseg2023, DialSeg711 dialseg711, TIAGE tiage, and MultiWOZ budzianowski2018multiwoz. Open-domain and semi-structured datasets include DailyDialog (synthetic concatenations) li2020dailydialog, Taskmaster byrne2019taskmaster, Topical-Chat topicalchat, and QMSum zhong2021qmsum. Several datasets provide sparse boundaries by design or annotate boundaries with intent (e.g., summarization units), which is precisely the setting where exact-match F1 becomes fragile. Together, these datasets span sparse, medium, and dense annotation regimes, providing a testbed for evaluating segmentation behavior under annotation granularity shift.
We refer to the SuperDialseg dataset as SuperSeg throughout. Separately, we denote the periodic oracle baseline family as Oracle-Periodic (Every-$N$).
Converting the datasets to a common canonical format is described in Appendix D.
@@ 5 Methodology
The experiments reported in Sections 6 and 7 intentionally evaluate only the neural boundary scoring component under static threshold++gap selection. Episodic, as deployed, is a hybrid system in which neural boundary scores are complemented by additional online signals, including embedding-drift–based change detection used by the commitment policy described in Appendix F. These auxiliary signals are not used in the main experiments and are held out deliberately to isolate the behavior of boundary-based evaluation metrics under controlled changes in boundary density. Accordingly, the performance tables and density–quality curves should be interpreted as diagnostics of scoring and selection behavior under controlled boundary densities.
@@ 5.1 Topic Segmentation Algorithm
Boundary density is controlled entirely by the boundary selection rule through the threshold $\tau$ and minimum spacing parameter $g$. In all experiments reported in this paper, $g$ is fixed to a single global value across all datasets, and $\tau$ is fixed globally after temperature calibration. Neither parameter is tuned per dataset, per split, or on held-out validation data to match gold boundary counts or to optimize any evaluation metric. As a result, variation in boundary density across datasets reflects differences in annotation granularity rather than dataset-specific hyperparameter adjustment. This choice is intentional: per-dataset tuning of $g$ or $\tau$ would partially obscure the granularity effects this work aims to expose.
@@
    Enforce a minimum spacing $g$ via greedy non-maximum suppression: candidates are processed in descending score order, and a candidate $i$ is accepted only if no previously accepted boundary lies within $g$ positions.
@@
Scope. For Figure 3, we allow the threshold $\tau$ to vary over time (logged as min_evidence, the running evidence threshold used by the selection rule); all other aspects of the selection rule remain unchanged. During evidence accumulation, candidate scores are evaluated against a fixed pre-change reference. Freezing both the pre-boundary context and the boundary-adjacent anchor turn ensures that the selection rule measures persistence of topic departure rather than repeated detection of the same topic transition.
@@
All main experimental results reported in Sections 6–7 use a fixed minimum spacing of $g=3$ messages and a fixed global threshold of $\tau=0.50$ applied to temperature-calibrated scores. These values are held constant across all datasets, splits, and runs.
@@ 5.3 Training Procedure
Stage 1: Synthetic Pretraining. We construct synthetic splice examples by concatenating real dialogue segments and labeling the splice point as a boundary. The synthetic splice corpus is constructed exclusively from training splits of the dialogue datasets listed in Table 1, ensuring no overlap with validation or test data. We remove boundary-local artifacts (e.g., greetings) to prevent spurious cues, and we evaluate a negative control set of single-segment examples to verify that the model does not hallucinate boundaries. While some synthetic splice points may join semantically related segments, this introduces conservative label noise that discourages hallucinated boundaries rather than inducing spurious segmentation.
@@
Importantly, temperature calibration rescales evidence but does not determine how many boundaries are produced. Boundary density is controlled exclusively by the selection rule (thresholding and spacing), not by temperature. While changing $T$ shifts the score distribution, it does not encode a target boundary rate and cannot, by itself, enforce consistent segmentation granularity across datasets. Calibration therefore improves the interpretability of scores but does not resolve granularity mismatch, which remains a decision-level issue handled by explicit boundary selection.
@@
Stage 2 Supervised fine-tuning SuperSeg, TIAGE Train only
@@
Table 1: Training, calibration, and evaluation protocol by stage. Stage 2 supervised fine-tuning uses only the training splits of SuperSeg and TIAGE to expose the scorer to both dense and sparse annotation regimes. Temperature calibration learns a single global temperature on a held-out validation split of SuperSeg. Stage 3 evaluation is performed on test splits only, including both datasets seen during training and datasets held out entirely from training and calibration.
@@ 6 Evaluation Metrics
We treat topic segmentation as predicting a set of boundary indices $P_{d}\subseteq\{1,\dots,T_{d}-1\}$ and compare against gold boundaries $G_{d}\subseteq\{1,\dots,T_{d}-1\}$. We compute precision $|P_{d}\cap G_{d}|/|P_{d}|$, recall $|P_{d}\cap G_{d}|/|G_{d}|$, and $F_{1}$ as the usual harmonic mean. For brevity, we refer to this metric simply as $F_{1}$ throughout.
@@ 6.1 Window-tolerant boundary accuracy (W-F1)
W-F1 measures boundary placement accuracy under a fixed tolerance window. A predicted boundary is counted as correct if it lies within a small window around a gold boundary, rather than requiring an exact index match. This makes W-F1 robust to small positional shifts that are often operationally irrelevant. The variant used here adopts a window-coverage convention, under which multiple predicted boundaries may fall within the tolerance window of a single gold boundary and be counted as correct. Formally, W-F1 is computed from dialogue-level precision and recall defined over boundary sets $P_{d}$ and $G_{d}$ using a tolerance window $\pm w$ (see Appendix A for the full definition).
@@
This window-coverage design is intentionally recall-favoring. Precision under W-F1 should therefore be interpreted as the fraction of predicted boundaries that fall near annotated topic transitions, rather than as a one-to-one measure of boundary set agreement. We adopt this convention to decouple boundary localization from boundary count, allowing W-F1 to serve as a diagnostic of whether boundaries are placed near annotated transitions. For this reason, W-F1 must be interpreted in conjunction with separate measures that account for boundary density and segment structure.
@@
Although our primary analysis uses many-to-one window-coverage matching, we additionally evaluate a standard one-to-one tolerant matching variant (maximum bipartite matching within the same tolerance window; see Appendix G to verify that the density-driven effects reported in this paper are not an artifact of the matching convention. Across both datasets, the resulting density–quality curves remain qualitatively unchanged (Figure 4.

### v3

@@ Abstract
Dialogue topic segmentation supports summarization, retrieval, memory management, and conversational continuity. Despite decades of prior work, evaluation practice remains framed as boundary prediction, using strict matching and F1-based metrics at a fixed threshold. This obscures boundary density and granularity effects—factors that matter as LLM-based conversational systems increasingly rely on segmentation to manage history beyond fixed context windows. Without structure, accumulated dialogue degrades efficiency and coherence.
This paper introduces an evaluation perspective for dialogue topic segmentation that reports boundary density and segment alignment diagnostics (purity and coverage) alongside window-tolerant F1 (W-F1). By explicitly analyzing boundary density through separating boundary scoring (per-position scores) from boundary selection (mapping scores to a boundary set), we evaluate segmentation quality across density regimes rather than at a single threshold setting. Through cross-dataset empirical evaluation, we show that reported performance differences across dialogue segmentation benchmarks are often not driven by boundary placement quality alone; under common evaluation protocols, annotation granularity mismatch and sparse boundary labels can dominate these differences.
We evaluate multiple segmentation strategies across eight dialogue datasets spanning task-oriented, open-domain, meeting-style, and synthetic interactions. Across these settings, boundary-based metrics are strongly coupled to boundary density: threshold sweeps produce larger W-F1 changes than switching between methods. These findings support viewing topic segmentation as a granularity selection problem and motivate treating boundary selection as a distinct, tunable step.
@@ 1 Introduction
The proposed separation of boundary scoring and boundary selection motivates alternative ways of reasoning about the construction, tuning, and evaluation of segmentation models under varying annotation granularities.
@@ 1.1 Motivation: Conversational Memory in AI Chat Systems
In this setting, topic segmentation functions as a memory control mechanism rather than a purely linguistic task. Poor segmentation can cause relevant context to be discarded prematurely or irrelevant context to be retained, leading to degraded long-horizon coherence and unstable behavior across turns, e.g., [sclar-etal-2024-quantifying, razavi-etal-2025-promptset]. The lack of robust long-horizon memory remains an open challenge for current LLM-based systems, particularly in conversations spanning multiple sessions, motivating careful attention to segmentation decisions and their evaluation.
@@ 2.1 Boundary-Based Evaluation Metrics
Topic segmentation evaluation has historically centered on boundary alignment. Early metrics such as $P_{k}$ beeferman1999statistical penalize segmentations where predicted and reference boundaries fall in different segments under a sliding window. WindowDiff pevzner2002critique refines this approach to address known biases in $P_{k}$. Both metrics are typically computed against a single reference segmentation and measure deviation from that reference, even though multiple segmentations (and granularities) can be defensible [fournier-inkpen-2012-segmentation, fournier-2013-evaluating]. More recent dialogue segmentation work has largely adopted precision, recall, and exact-match $F_{1}$ computed over boundary positions [jiang2023superdialseg, xia2022dialogue]. These boundary-only scores likewise treat the available annotation as the evaluation target; when annotation granularity varies across datasets or differs from system behavior, they can yield misleading comparisons.
@@
A substantial line of work in text segmentation evaluation argues that (1) strict boundary matching over-penalizes near misses, and (2) window-based metrics such as $P_{k}$ and WindowDiff introduce task- and parameter-dependent biases, motivating edit-distance and confusion-matrix-based alternatives. fournier-inkpen-2012-segmentation propose segmentation similarity ($S$), an edit-distance–based similarity that is symmetric (does not privilege a single “true” reference) and supports evaluation against multiply-coded corpora. fournier-2013-evaluating extends this direction with boundary edit distance and boundary similarity ($B$), yielding a principled way to obtain boundary-level IR-style quantities (precision/recall/$F$) while awarding graded credit for near misses. Relatedly, scaiano-inkpen-2012-getting derive a boundary confusion matrix from WindowDiff-style windows (WinPR), separating false positives from false negatives while retaining tunable near-miss sensitivity. For corpora with multiple annotators or where segmentation is treated as a unitizing task, agreement measures that jointly align and score units (e.g., $\gamma$) provide an alternative to single-reference evaluation [mathet2015gamma].
@@ 2.2 Dialogue Segmentation Datasets
Dialogue segmentation benchmarks differ not only in domain but in what counts as a topic boundary: some datasets provide explicit segment boundaries, others provide weaker supervisory signals (e.g., section/domain labels) that must be repurposed, and some evaluation sets are synthetically constructed via concatenation. These design choices induce large, systematic differences in boundary density and implied granularity, so boundary-only metrics can confound localization with density alignment. Section 4 summarizes the eight datasets used in our experiments and the exact gold-boundary definitions used after canonicalization.
@@ 4 Datasets
We evaluate across eight dialogue datasets spanning diverse annotation regimes (see Section 2.2 for background on annotation philosophies). The purpose of cross-dataset evaluation is not only to compare performance, but to expose how annotation philosophy and boundary density interact with evaluation metrics. We summarize the supervision signal used for each dataset and how it is converted to turn-adjacent gold boundaries in Table 1 (full canonicalization details in Appendix D).
Explicit segmentation annotation
DialSeg711 xu2021topicaware is constructed by joining single-topic dialogues (from MultiWOZ and the Stanford Dialogue Dataset), so boundaries primarily reflect splice points rather than naturally occurring gradual topic transitions. SuperDialseg jiang2023superdialseg provides large-scale supervised segmentation built from document-grounded dialogues, with segment IDs aligned to source document structure. TIAGE xie2021tiage augments PersonaChat with per-turn topic-shift annotations, yielding sparse shift points rather than hierarchical segment structure. QMSum zhong2021qmsum provides topic annotations for meeting transcripts; because QMSum allows topic-relevant spans to be non-contiguous, we convert its annotations to turn-adjacent boundary positions as described in Appendix D.
Repurposed labels
MultiWOZ budzianowski2018multiwoz provides domain and dialogue-state annotations; we derive boundaries from domain label changes. DailyDialog li2017dailydialog, Taskmaster byrne2019taskmaster, and Topical-Chat gopalakrishnan2019topicalchat do not provide explicit segmentation boundaries; we derive boundaries from changes in available topic/domain/task labels between consecutive utterances (Appendix D).
Together, these datasets span sparse, medium, and dense annotation regimes, providing a testbed for evaluating segmentation behavior under annotation granularity shift. We refer to SuperDialseg as SuperSeg throughout.
Dataset Supervision signal Gold boundary derivation
Explicit segmentation / shift labels
DialSeg711 synthetic splice construction splice point (segment ID change)
SuperSeg document section IDs section change
TIAGE topic-shift annotations annotated shift / segment change
QMSum topic spans boundary at adjacent span break
Repurposed labels (boundary = label change)
MultiWOZ domain label domain change
DailyDialog dialogue topic label topic change
Taskmaster task/subtask labels task/subtask change
Topical-Chat conversation topic label topic change
Table 1: Supervision signals and the resulting turn-adjacent gold boundaries used in evaluation. Datasets are cited in the main text (§4); full extraction/canonicalization is in Appendix D.
@@ 5 Methodology
The experiments reported in Sections 6 and 7 intentionally evaluate only the neural boundary scoring component under static threshold++gap selection. Episodic, as deployed, is a hybrid system in which neural boundary scores are complemented by additional online signals (e.g., embedding-drift–based change detection). These auxiliary signals are not used in the main experiments and are held out deliberately to isolate the behavior of boundary-based evaluation metrics under controlled changes in boundary density. Accordingly, the performance tables and density–quality curves should be interpreted as diagnostics of scoring and selection behavior under controlled boundary densities.
@@ 5.1 Topic Segmentation Algorithm
Boundary density is controlled entirely by the boundary selection rule through the threshold $\tau$ and minimum spacing parameter $g$. For all single-operating-point results reported in this paper (Tables 4, 5, 10, and 11), $g$ is fixed at 3 and $\tau$ is fixed at 0.50 after temperature calibration. Density–quality curves (Section 7.1) sweep $\tau$ while holding $g$ fixed; the adaptive controller (Section 9.1, Figure 3) is a separate control experiment. Neither parameter is tuned per dataset, per split, or on held-out validation data to match gold boundary counts or to optimize any evaluation metric. As a result, variation in boundary density across datasets reflects differences in annotation granularity rather than dataset-specific hyperparameter adjustment. This choice is intentional: per-dataset tuning of $g$ or $\tau$ would partially obscure the granularity effects this work aims to expose.
@@
    Enforce a minimum spacing $g$ via greedy non-maximum suppression: candidates are processed in descending score order, and a candidate i is accepted only if $|i-b|\geq g$ for all previously accepted boundaries $b$.
@@
Scope. For Figure 3, we allow the threshold $\tau$ to vary over time (logged as min_evidence, the running evidence threshold used by the selection rule); all other aspects of the selection rule remain unchanged. During evidence accumulation, each candidate’s fixed score $s_{i}$ (computed once at first availability using frozen reference windows) is repeatedly added to its evidence total, implementing a time-to-threshold integrator: candidates with stronger initial boundary signals reach the commitment threshold faster. Freezing the score at the point of first availability ensures that the commitment decision reflects the strength of the boundary signal at the moment of topic departure, rather than being confounded by subsequent context that may diverge further or revisit the original topic.
@@
All single-operating-point results for the proposed neural scorer (Tables 10 and 11) use a fixed minimum spacing of $g=3$ messages and a fixed global threshold of $\tau=0.50$ applied to temperature-calibrated scores. These values are held constant across all datasets, splits, and runs. Baselines in Table 4 that directly output boundaries (e.g., No-boundary, Periodic, Oracle-density Random) do not use $\tau$/$g$; external methods in Table 5 (TextTiling, CSM) use their own internal scoring and thresholding mechanisms evaluated at fixed operating points (see Appendix H for details). Density–quality curves (Section 7.1, Figure 1) sweep $\tau$ over $[0.05,0.95]$ while holding $g=3$ fixed, as described in “Threshold sweeps for density–quality curves” above.
@@ 5.3 Training Procedure
Stage 1: Synthetic Pretraining. We construct synthetic splice examples by concatenating real dialogue segments and labeling the splice point as a boundary. The synthetic splice corpus is constructed from training splits of DailyDialog and MultiWOZ, ensuring no overlap with validation or test data. We remove boundary-local artifacts (e.g., greetings) to prevent spurious cues, and we evaluate a negative control set of single-segment examples to verify that the model does not hallucinate boundaries. While some synthetic splice points may join semantically related segments, this introduces conservative label noise that discourages hallucinated boundaries rather than inducing spurious segmentation.
@@
Importantly, temperature calibration does not explicitly enforce a target boundary rate. Because it rescales scores, changing $T$ can affect boundary density under a fixed threshold; however, explicit density control requires selection parameters ($\tau$, $g$) or a controller. Temperature calibration alone cannot enforce consistent segmentation granularity across datasets. Calibration therefore improves the interpretability of scores but does not resolve granularity mismatch, which remains a decision-level issue handled by explicit boundary selection.
@@
Stage 2 Supervised fine-tuning SuperSeg, TIAGE Train (val for metrics)
@@
Table 2: Training, calibration, and evaluation protocol by stage. Stage 2 supervised fine-tuning uses only the training splits of SuperSeg and TIAGE to expose the scorer to both dense and sparse annotation regimes. Temperature calibration learns a single global temperature on a held-out validation split of SuperSeg. Stage 3 evaluation is performed on test splits only, including datasets seen during supervised fine-tuning (SuperSeg, TIAGE), datasets used only for Stage 1 synthetic pretraining (DailyDialog, MultiWOZ), and datasets held out entirely (DialSeg711, Taskmaster, Topical-Chat, QMSum).
@@ 6 Evaluation Metrics
We treat topic segmentation as predicting a set of boundary indices $P_{d}\subseteq\{1,\dots,T_{d}-1\}$ and compare against gold boundaries $G_{d}\subseteq\{1,\dots,T_{d}-1\}$. We compute precision $|P_{d}\cap G_{d}|/|P_{d}|$, recall $|P_{d}\cap G_{d}|/|G_{d}|$, and $F_{1}$ as the usual harmonic mean. For brevity, we refer to this metric simply as $F_{1}$ throughout. Dataset-level F1 is micro-averaged: $\mathrm{F1}=2\sum_{d}|P_{d}\cap G_{d}|\;/\;(\sum_{d}|P_{d}|+\sum_{d}|G_{d}|)$.
@@ 6.1 Window-tolerant boundary accuracy (W-F1)
W-F1 measures boundary placement accuracy under a fixed tolerance window. A predicted boundary is counted as correct if it lies within a small window around a gold boundary, rather than requiring an exact index match. This makes W-F1 robust to small positional shifts that are often operationally irrelevant. The variant used here adopts a window-coverage convention, under which multiple predicted boundaries may fall within the tolerance window of a single gold boundary and be counted as correct. Formally, W-F1 is computed from dialogue-level precision and recall defined over boundary sets $P_{d}$ and $G_{d}$ using a tolerance window $\pm w$ (see Appendix A for the full definition). In all experiments here, we use $w=1$.
@@
This window-coverage design is intentionally duplicate-tolerant: multiple predictions near the same gold boundary are not penalized as false positives. Precision under W-F1 should therefore be interpreted as the fraction of predicted boundaries that fall near annotated topic transitions, rather than as a one-to-one measure of boundary set agreement. We adopt this convention to decouple boundary localization from boundary count, allowing W-F1 to serve as a diagnostic of whether boundaries are placed near annotated transitions. For this reason, W-F1 must be interpreted in conjunction with separate measures that account for boundary density and segment structure.
@@
Although our primary analysis uses many-to-one window-coverage matching, we additionally evaluate a standard one-to-one tolerant matching variant (maximum bipartite matching within the same tolerance window); see Appendix G to verify that the density-driven effects reported in this paper are not an artifact of the matching convention. Across both datasets, the resulting density–quality curves remain qualitatively unchanged (Figure G.1).